# Can industrial overcapacity enable seasonal flexibility in electricity use? A case study of aluminum smelting in China

### v1

@@ 4.2 Aluminum demand projections
Our objective is to project China’s primary aluminum demand for the coming decades, which requires projecting both total aluminum demand and the future supply of recycled aluminum. Our methodology is based on the assumption that per capita aluminum in-use stock tends to saturate as the economy develops, a common approach for projecting the demand for metal raw materials like steel and aluminum [song_mapping_2020, li_analysis_2022]. To obtain historical aluminum in-use stock data, which is not explicitly tracked in official statistics, we sourced historical aluminum production data from the China Statistical Yearbook. We then calculated the aluminum stock by accumulating historical production and accounting for product retirement. Specifically, using China’s primary and secondary aluminum production data from 1993 to 2023 and assuming that the lifespan of aluminum products follows a Weibull distribution, we calculated the annual aluminum in-use stock up to 2024. Next, we input this data into the EnergyPathways model [Farid2024], a mature, open-source software for forecasting various types of demand, which has been used in the Net-Zero America [larson2020net] and Net-Zero Australia [mccall2025topical] projects. Using the stock-based method, with the projected population by chen_provincial_2020 as the driver, we selected logistic regression as the extrapolation curve for future aluminum demand. The saturation level is set differently for various scenarios, as detailed in Supplementary Note 1. The EP model projects the annual aluminum stock, total aluminum demand, and retired aluminum products (the available scrap in that year) up to 2060. The calculation formula is as follows:
@@
Here, $\text{Stock}_{\text{year}}$ represents the extrapolated aluminum stock, $\text{Production}_{\text{year}}$ is the total demand/production for aluminum, and scrap is calculated from the historical stock based on the assumption that product lifespan follows a Weibull distribution. Therefore, the total aluminum demand and available scrap can be calculated from the above formula. Based on the scrap, we can calculate the recycled aluminum production ($\text{Production}_{\text{recycle}}=\text{Scrap}\times\text{Recycle}_{\text{rate}}$). By subtracting the recycled aluminum production from the total aluminum demand for that year, we obtain the primary aluminum demand/production ($\text{Production}_{\text{primary}}=\text{Production}-\text{Production}_{\text{recycle}}$). China is also a major aluminum exporter. Based on current export levels, we have also established two scenarios: reduced exports and increased exports (see Table 1), with further details provided in Supplementary Note 1. Having obtained the annual aluminum demand projections, we assume that the hourly aluminum demand remains constant throughout the year, which simplifies demand distribution yet imposes a stricter supply-demand balance. This demand must be satisfied either through production from aluminum smelters or by drawing from existing aluminum inventories.
@@ 4.3 Modeling smelter operational flexibility
This study employs a mixed-integer linear model, an initial version of which was developed by Shen et al. [shen_improved_2025] as a linear model for production and energy consumption decisions in smelter operations. The model is capable of representing the relationships between internal temperature changes and electrolysis rates in response to applied voltage and current in a potline, as observed in both numerical simulations and experiments [oye_power_2011, oye_power_nodate]. It is transformed into a linear programming model through piecewise linearization and numerical approximation. To model the shutdown and startup decisions and their associated costs for smelters, we introduced startup and shutdown constraints and costs into the aforementioned linear model, transforming it into a mixed-integer model (see Supplementary note 3 and 5). This enables our model to optimize the smelter’s operational decisions over a longer timescale. However, the introduction of integer variables makes it computationally infeasible to directly embed a plant-level operational model into a power system planning model. Therefore, we design a decomposition-based iterative algorithm for solving this problem. To determine the production capacity of smelters, based on the Chinese government’s policy that primary aluminum capacity will not exceed 45 million tonnes per year, we assume that the production capacity of aluminum smelters remains unchanged at the 2024 level of 45 million tonnes per year, unless early retirement is chosen. This assumption is based on the typical lifespan of an aluminum smelter, which exceeds 50 years, and the fact that more than 90% of China’s aluminum smelters were built after 2000.
Since not all provinces have complete statistical data on aluminum smelting capacity, we distributed the national capacity to each province based on its 2023 production. We assumed that the capacity share of each province remains unchanged when reducing overcapacity (see Supplementary Note 3.3 for more details). This approach is justified by the fact that China’s aluminum smelting industry currently operates at a high capacity utilization rate of approximately 97%, meaning that most provinces are running at near full capacity. For estimating the needed workforce for smelters and power plants, we use the average number of workers required per unit of operating capacity, assuming a certain proportion of variable workforce (see Supplementary Note 3.5).
@@ 4.4 Smelter costing
The investment, operation, and maintenance costs of aluminum smelters are estimated from annual reports from Chinese smelters [CHALCO2020AnnualReport]. However, it is challenging to determine the idling and startup/shutdown costs, as there are not many historical cases where smelters were idled for long periods or shut down due to insufficient demand. According to publicly available reports, the normalized cost for restarting potlines can vary from 110,000 CNY/MW to 760,000 CNY/MW [Hu2009, noauthor_high_2020, noauthor_alcoa_2021], which is 1000 times higher than restarting coal-fired units [smith_analysis_2019]. This study does not attempt to predict the future trajectory of smelter restart costs, but rather presents extreme cases in addition to the baseline case to represent the two extremes of the spectrum for the development of flexible aluminum smelting technologies, resulting in a total of four smelter flexibility cases. In the Low-flexibility and Mid-flexibility cases, the smelter flexibility parameters are set based on current literature and reports. The former’s restart cost is estimated at 760,000 CNY/MW in a case of a U.S.-owned smelter. The latter’s restart cost is from a documented restart event at a smelter in Yunnan, China, estimated at 110,000 CNY/MW. For smelters to conduct seasonal production and help adapt to fluctuations in the power grid’s supply and demand, they may need to increase the storage period of aluminum products to act as a buffer for balancing production and demand. Therefore, we also need to model the warehouse costs of aluminum products. Product warehouse costs are estimated by multiplying the actual short-term warehouse rental fees in each region for 2023 by the required warehouse area [Colliers2024Logistics]. However, this may lead to an overestimation, as long-term rental fees are typically lower. A more detailed description of the smelter operational cost calculation method is in Supplementary note 3.3.
@@ 4.5 Power system capacity expansion model
This study aims to evaluate the load flexibility value of aluminum smelters for China’s decarbonized power system. A provincial resolution is needed because China’s aluminum industry is unevenly distributed across multiple provinces, and the highly interconnected nature of China’s power grid means that the impact of this flexibility is likely to span across provincial boundaries. We set the net-zero year of 2050, even though China’s overall carbon neutrality target is 2060 [yang_chinas_2024], since the power sector is generally considered the primary and critical area for achieving this goal and requires deep decarbonization at an earlier stage. According to several studies and scenario analyses, China’s power sector is expected to achieve net-zero emissions between 2045 and 2055 [liu_chinas_2025]. To compare the effects of different carbon reduction constraints, we set up scenarios with 20%, 60%, and 100% reductions in carbon emissions for 2030, 2040, and 2050, respectively, based on 2020 emission levels.
This study uses PyPSA, an open-source model described in detail elsewhere [PyPSA], to explore China’s pathway to carbon emission reduction. The model uses 2020 as the base year to formulate a brownfield capacity expansion model, covering thirty-one provincial administrative regions of China, with each province aggregated into a single node. Each node not only handles electricity but also incorporates multiple energy carriers, including heat, hydrogen, biomass, methane, and coal, reflecting the integration of a multi-energy system. The connections between nodes include electricity transmission via high-voltage direct current (HVDC) and high-voltage alternating current (HVAC) lines, as well as hydrogen transmission via hydrogen pipelines. The PyPSA model uses linear optimization to minimize annual operational and investment costs while adhering to technical and physical constraints. The system operation is modeled with an 8,760-hour time resolution over a single meteorological year (2020 as the representative year). The optimization includes the capacity of generators, storage, converters, and transmission, as well as the hourly operational dispatch of each technology. Exogenous variables include: demand curves for different sectors, the ratio of district heating to decentralized heating, the capacity of hydro reservoirs and run-of-river generators, and pumped hydro storage capacity. HVAC and HVDC lines are assumed to be long-lasting and therefore permanent. Thus, transmission network assets can be reinforced but not removed. The detailed input data for the model are primarily derived from the PyPSA China-specific project, originally established and open-sourced by Zhou et al. [zhou_multienergy_2024], and have been used in peer-reviewed publications.
@@ 4.6 Dimensionality reduction and tailored iterative solution
Our objective is to address the computational intractability that arises when integrating detailed smelter operation models (which include integer variables) into a large-scale power system planning model. We adopt an iterative decomposition framework based on the D3R method proposed by Lyu et al. [lyu_data-driven_2025], which provides a computationally efficient linear approximation of the complex industrial load constraints. The process begins by embedding this linearized smelter model obtained through the D3R framework into the main linear capacity expansion model. Solving this initial integrated model yields a first estimate of electricity system capacity expansion results and the hourly local marginal prices (LMPs) for electricity in each province. Then, these LMPs are fixed and used as an exogenous price signal for the detailed smelter operational models. This step, a core part of the decomposition, allows each smelter’s 8760-hourly cost-minimization problem (including all unit commitment constraints) to be solved independently and in parallel. Once the detailed optimal operational schedule (i.e., the 8760-hourly load profile) is determined for each smelter, we fix these schedules in the main capacity expansion model and then re-optimize all remaining power system planning and operational variables. This re-optimization generates a new, updated set of LMPs, which are fed back into the detailed smelter models. This entire process is iterated until the objective function of the planning model converges—which our tests show occurs within several iterations—indicating a consistent, optimal solution for both system planning and smelter operation has been found. For more details, see Supplementary note 5.
@@ 4.7 Limitations and opportunities
Our study has several limitations that offer avenues for future research. Our model employs a top-down approach, projecting nationwide aluminum demand, which may differ from granular end-use forecasts [lennon_aluminium_2022]. However, our projected trends align with those of bottom-up studies [yang_supply_2024]. We also simplify the distinction between primary/recycled aluminum, as well as market/trade dynamics. Regarding smelter representation, our model aggregates plants at the provincial level (see Supplementary Note 5.2 for details) and, due to computational constraints, assumes fixed capacity ratios between provinces. Future work could leverage efficient frameworks to optimize the spatial distribution of smelter retirement or migration, thereby capturing regional flexibility value, or, with enough data, offer plant-level operational strategies. We also do not differentiate between smelting technologies, retrofitting options, or stoichiometry of aluminum smelting [shen_technologies_2024, tan_different_2025]. Finally, our co-optimization focuses on the aluminum smelting industry — the most electrified energy-intensive industry and one that is roughly ten times more energy-intensive per tonne of product than steelmaking [IAI2023Smelting, Hasanbeigi2019]. The aluminum smelting’s high electricity use and operational flexibility make it a prime candidate for studying grid interaction, but subsequent research should explore the co-optimization of multiple energy-intensive sectors to develop a more general understanding of how other energy-intensive industries can collectively manage structural overcapacity and navigate the low-carbon transition. More broadly, the question of how to leverage industrial-energy sector coupling to mitigate the socio-economic impacts of energy decarbonization and industrial restructuring represents a promising future research direction.
@@ Acknowledgment
The authors would like to thank Eric Larson for reviewing the paper, Jiaxiang Zhang, Xicheng Cha, Jiarui Li, and Yingrui Zhang for their assistance with data collection, and Zongzhi Du for providing consultation on the operation of aluminum corporations in China.

### v2

@@ 4.2 Aluminum demand projections
Our objective is to project China’s primary aluminum demand for the coming decades, which requires projecting both total aluminum demand and the future supply of recycled aluminum. Our methodology is based on the assumption that per capita aluminum in-use stock tends to saturate as the economy develops, a common approach for projecting the demand for metal raw materials like steel and aluminum [51, 31]. To obtain historical aluminum in-use stock data, which is not explicitly tracked in official statistics, we sourced historical aluminum production data from the China Statistical Yearbook. We then calculated the aluminum stock by accumulating historical production and accounting for product retirement. Specifically, using China’s primary and secondary aluminum production data from 1993 to 2023 and assuming that the lifespan of aluminum products follows a Weibull distribution, we calculated the annual aluminum in-use stock up to 2024. Next, we input this data into the EnergyPathways (EP) model [17], a mature, open-source software for forecasting commodity demands in energy systems, which has been used in high-profile studies such as the Net-Zero America [28] and Net-Zero Australia [13] projects. The choice of EP is motivated by its proven capability to model future product demand. Using the stock-based saturation method and projected future population from Chen et al. [7], we used logistic regression (S-curve fitting) to extrapolate future aluminum demand. This approach is well-suited for our study as it captures the decline in primary metal consumption due to slowing population growth, despite continuing economic growth—a critical feature for analyzing the upcoming overcapacity in China’s aluminum sector as it matures. The EP model projects the annual aluminum stock, total aluminum demand, and retired aluminum products (the available scrap in that year) up to 2060. The calculation formula is as follows:
@@
Here, $\text{Stock}_{\text{year}}$ represents the extrapolated aluminum stock, $\text{Production}_{\text{year}}$ is the total demand/production for aluminum, and scrap is calculated from the historical stock based on the assumption that product lifespan follows a Weibull distribution. Therefore, the total aluminum demand and available scrap can be calculated from the above formula. Based on the scrap, we can calculate the recycled aluminum production ($\text{Production}_{\text{recycle}}=\text{Scrap}\times\text{Recycle}_{\text{rate}}$). By subtracting the recycled aluminum production from the total aluminum demand for that year, we obtain the primary aluminum demand/production ($\text{Production}_{\text{primary}}=\text{Production}-\text{Production}_{\text{recycle}}$). China is also a major aluminum exporter. Based on current export levels, we have also established two scenarios: reduced exports and increased exports (see Table 1), with further details provided in Supplementary Note 2.4. Having obtained the annual aluminum demand projections, we assume that the hourly aluminum demand remains constant throughout the year, which simplifies demand distribution yet imposes a stricter supply-demand balance. This demand must be satisfied either through production from aluminum smelters or by drawing from existing aluminum inventories.
@@ 4.3 Modeling smelter operational flexibility
This study employs a mixed-integer linear model, an initial version of which was developed by Shen et al. [47] as a linear model for production and energy consumption decisions in smelter operations. The model is capable of representing the relationships between internal temperature changes and electrolysis rates in response to applied voltage and current in a potline, as observed in both numerical simulations and experiments [41, 42]. It is transformed into a linear programming model through piecewise linearization and numerical approximation. To capture the unique flexibility characteristics of aluminum smelters, we introduce binary variables for startup and shutdown decisions, which extends this framework to a mixed-integer linear programming (MILP) model. This extension is essential because aluminum smelters, similar to thermal power units, exhibit discrete operational states and path-dependent costs. Specifically, if a potline remains operational, it must maintain a minimum power level (typically 70–90% of rated power) to preserve the internal thermal balance of the electrolytic cells. Reducing power further requires a complete shutdown, which then incurs significant restart costs to restore the thermal and chemical equilibrium—costs that can be orders of magnitude higher than those for conventional power units [48]. A pure linear model would fail to capture these ”all-or-nothing” and ”minimum-stable-output” constraints, potentially overestimating the smelter’s ability to provide continuous, low-cost load shedding. This enables our model to optimize the smelter’s operational decisions over a longer timescale (see Supplementary Note 3 and 5). However, the introduction of integer variables makes it computationally infeasible to directly embed a plant-level operational model into a power system planning model. Therefore, we design a decomposition-based iterative algorithm for solving this problem. To determine the production capacity of smelters, based on the Chinese government’s policy that primary aluminum capacity will not exceed 45 million tonnes per year, we assume that the production capacity of aluminum smelters remains unchanged at the 2024 level of 45 million tonnes per year, unless early retirement is chosen. This assumption is based on the typical lifespan of an aluminum smelter, which exceeds 50 years, and the fact that more than 90% of China’s aluminum smelters were built after 2000 (see Supplementary Note 3.3).
Since not all provinces have complete statistical data on aluminum smelting capacity, we distributed the national capacity to each province based on its 2023 production. We assumed that the capacity share of each province remains unchanged when reducing overcapacity. This approach is justified by the fact that China’s aluminum smelting industry currently operates at a high capacity utilization rate of approximately 97%, meaning that most provinces are running at near full capacity. For estimating the needed workforce for smelters and power plants, we use the average number of workers required per unit of operating capacity, assuming a certain proportion of variable workforce (see Supplementary Note 3.7).
@@ 4.4 Smelter costing
The investment, operation, and maintenance costs of aluminum smelters are estimated from annual reports from Chinese smelters [2]. However, it is challenging to determine the idling and startup/shutdown costs, as there are not many historical cases where smelters were idled for long periods or shut down due to insufficient demand. According to publicly available reports, the normalized cost for restarting potlines can vary from 110,000 CNY/MW to 760,000 CNY/MW [60, 29, 11], which is 1000 times higher than restarting coal-fired units [48]. This study does not attempt to predict the future trajectory of smelter restart costs, but rather presents extreme cases in addition to the baseline case to represent the two extremes of the spectrum for the development of flexible aluminum smelting technologies, resulting in a total of four smelter flexibility cases. In the Low-flexibility and Mid-flexibility cases, the smelter flexibility parameters are set based on current literature and reports. The former’s restart cost is estimated at 760,000 CNY/MW in a case of a U.S.-owned smelter. The latter’s restart cost is from a documented restart event at a smelter in Yunnan, China, estimated at 110,000 CNY/MW. For smelters to conduct seasonal production and help adapt to fluctuations in the power grid’s supply and demand, they may need to increase the storage period of aluminum products to act as a buffer for balancing production and demand. Therefore, we also need to model the warehouse costs of aluminum products. Product warehouse costs are estimated by multiplying the actual short-term warehouse rental fees in each region for 2023 by the required warehouse area [10]. However, this may lead to an overestimation, as long-term rental fees are typically lower. A more detailed description of the smelter operational cost calculation method is in Supplementary Note 3.3.
@@ 4.5 Power system capacity expansion model
This study evaluates the value of industrial load flexibility within a national-scale, multi-period capacity expansion framework. A provincial resolution is needed because China’s aluminum industry is unevenly distributed across multiple provinces, and the highly interconnected nature of China’s power grid means that the impact of this flexibility is likely to span across provincial boundaries. The modeling aims to identify the least-cost infrastructure investment and operational dispatch decisions needed to meet climate targets. While China’s national target is to achieve carbon neutrality by 2060 [61], we define a sector-specific 2050 net-zero scenario for the power and heating sectors (DP-2050). This is because easier-to-abate sectors may be decarbonized earlier in order to achieve economy-wide net-zero emissions by 2060. Power and residential heating are relatively easy-to-abate sectors due to the rapid decline in renewable energy costs and the maturity of technologies such as heat pumps, and thus achieving net-zero in power and heating sectors by 2050 may be a necessary prerequisite for the decarbonization of other end-use sectors. Furthermore, according to several studies and scenario analyses, China’s power sector is expected to reach net-zero emissions between 2045 and 2055 [35], allowing the sector to transition toward net-negative emissions (e.g., via bioenergy with carbon capture and storage) to offset residual emissions from hard-to-abate sectors by 2060. This timeline is supported by studies from the International Energy Agency [25], Tsinghua University [23], and the Rocky Mountain Institute [43]. To compare the effects of different carbon reduction constraints, we set up scenarios with 20%, 60%, and 100% reductions in carbon emissions for 2030, 2040, and 2050, respectively, based on 2020 emission levels.
We utilize the Python for Power System Analysis (PyPSA) framework, a widely-recognized, open-source tool for high-resolution energy system optimization. PyPSA has been extensively applied in numerous peer-reviewed studies to explore large-scale decarbonization pathways, most notably within the European context through the PyPSA-Eur project [22]. Unlike “copper plate” models, PyPSA captures the essential physics of transmission through a linearized DC approximation of power flow, enforcing Kirchhoff’s laws and identifying network bottlenecks that drive storage requirements and renewable curtailment. Furthermore, PyPSA natively handles sector coupling, allowing electricity, heat, hydrogen, and gas to be co-optimized—a requirement for capturing the interaction between industrial electrification and electricity sector dynamics. In this study, we utilize PyPSA-China, a peer-reviewed model build on PyPSA that incorporates the specialized dataset for China’s net-zero transition analysis [65].
@@ 4.6 Dimensionality reduction and tailored iterative solution
Our objective is to address the computational intractability that arises when integrating detailed smelter operation models (which include integer variables) into a large-scale power system planning model. Embedding millions of system variables alongside thousands of discrete industrial constraints would result in a “curse of dimensionality” that is numerically unsolvable. We adopt an iterative decomposition framework based on the Data-Driven Dimension Reduction (D3R) method proposed by Lyu et al. [37]. Unlike traditional analytical approximations that rely on pre-specified geometric templates (e.g., box or ellipsoid constraints), D3R utilizes inverse optimization to “learn” the feasible operating region of complex industrial loads. It trains a low-dimensional polytope that best approximates the smelter’s physical flexibility (including thermal and stoichiometry constraints) while remaining compatible with linear optimization. The D3R framework has been rigorously verified for its ability to maintain high approximation accuracy while reducing the computational burden by orders of magnitude, making it uniquely suitable for modeling the flexibility potential of energy-intensive industries in large-scale energy systems [37].
The process begins by embedding this linearized smelter model obtained through the D3R framework into the main linear capacity expansion model. Solving this initial integrated model yields a first estimate of electricity system capacity expansion results and the hourly local marginal prices (LMPs) for electricity in each province. Then, these LMPs are fixed and used as an exogenous price signal for the detailed smelter operational models. This step, a core part of the decomposition, allows each smelter’s 8760-hourly cost-minimization problem (including all unit commitment constraints) to be solved independently and in parallel. Once the detailed optimal operational schedule (i.e., the 8760-hourly load profile) is determined for each smelter, we fix these schedules in the main capacity expansion model and then re-optimize all remaining power system planning and operational variables. This re-optimization generates a new, updated set of LMPs, which are fed back into the detailed smelter models. This entire process is iterated until the objective function of the planning model converges—which our tests show occurs within several iterations—indicating a consistent, optimal solution for both system planning and smelter operation has been found. For more details, see Supplementary Note 5.
@@ 4.7 Limitations and opportunities
Our study has several limitations that offer avenues for future research. Our model employs a top-down approach, projecting nationwide aluminum demand, which may differ from granular end-use forecasts [30]. However, our projected trends align with those of bottom-up studies [62]. We also simplify the distinction between primary/recycled aluminum, as well as market/trade dynamics. Regarding smelter representation, our model aggregates plants at the provincial level (see Supplementary Note 5.2 for details) and, due to computational constraints, assumes fixed capacity ratios between provinces. This spatial aggregation likely results in a conservative estimate of flexibility value, as it ignores facility-level modularity—in which individual potlines or plants could be shut down more efficiently than provincial-level ramping—and overlooks intra-provincial grid constraints and technological heterogeneity among plants. The fixed capacity ratio assumption further prevents us from deriving an optimized, facility-level retirement strategy. In reality, a more granular representation of retirement decisions would likely favor retaining plants that are more modern and technically efficient, which would be better equipped to provide grid flexibility. Future work could leverage efficient frameworks to optimize the spatial distribution of smelter retirement or migration, thereby capturing regional flexibility value, or, with enough data, offer plant-level operational strategies. We also do not differentiate between smelting technologies, retrofitting options, or stoichiometry of aluminum smelting [46, 53]. Primary aluminum is uniquely suited for seasonal stockpiling due to its homogeneity and negligible storage degradation, whereas other commodities (e.g., steel, cement) face higher physical risks and holding costs. When extending this framework to other heavy industries, future research must incorporate sector-specific storage degradation and holding costs. Finally, our co-optimization focuses on the aluminum smelting industry — the most electrified energy-intensive industry. Subsequent research should explore the co-optimization of multiple energy-intensive sectors to develop a more general understanding of how other energy-intensive industries can collectively manage structural overcapacity and navigate the low-carbon transition. More broadly, the question of how to leverage industrial-energy sector coupling to mitigate the socio-economic impacts of energy decarbonization and industrial restructuring represents a promising future research direction.
@@ Acknowledgment
The authors thank Tengmu Li for his valuable perspectives on the seasonal socio-economic rhythms of the contemporary workforce, Zongzhi Du for providing consultation on the operation of aluminum corporations in China, and Eric Larson for reviewing the paper.